# Constant-Memory Strategies in Stochastic Games: Best Responses and Equilibria

### v1

@@
Abstract
Stochastic games have become a prevalent framework for studying long-term multi-agent interactions, especially in the domain of multi-agent reinforcement learning. However, the notion of equilibria in a stochastic game is fraught with subtleties, mainly because a strategy in a stochastic game can potentially map any observed history to (possibly randomized) actions, leading to an infinitely large strategy space for each agent. In particular, it has been proven that there exists a Nash equilibrium implemented by a profile of stationary strategies. This existence result has largely drawn the community’s attention to devising multi-agent learning algorithms to reach a stationary equilibrium. Nevertheless, in real-world scenarios, particularly when one has only limited knowledge of her opponents, the assumption that all opponents will use stationary strategies is indeed too strong. To this end, in this paper, we comprehensively investigate a less restricted class of strategies called constant-memory strategies, which map constant-length history segments to actions. Despite being weaker than Turing-machine strategies, constant-memory strategies can lead to better outcomes than those under stationary strategies. We show that given a constant-memory strategy, there always exists a deterministic constant-memory strategy that uses the same length of memory as a pure strategy best response. In addition, we provide an existence result for equilibria, indicating that given any finite length of memory, there always exists a Nash equilibrium in which all agents adopt constant-memory strategies using that length of memory. However, we also show that the best response against a mixed constant-memory strategy, as if an agent is in a tournament with known opponent types, is not constant-memory and is even hard to compute. Finally, based on the aforementioned theoretical takeaways, we present simple parameterized methods to generate single-agent RL environments that can be arbitrarily difficult to learn, achievable through slight modifications of the original multi-agent environment. We hope this work significantly deepens the understanding of theoretical issues in single-agent planning under multi-agent systems and uncovers the connection between decision models in single-agent scenarios and those in multi-agent scenarios.
@@ Abstract
Keywords: Stochastic games; Bounded rationality; Best response; Restricted memory; Reinforcement learning
1 Introduction
@@
Various real-world situations that involve long-term interactions among a group of participants can be modeled as stochastic games, such as negotiation between multiple stakeholders [8, 17, 7], bidding and mechanism design in repeated auctions [25, 9, 22, 21, 47, 43], multi-agent teamwork [42, 34, 46], and even human-robot collaboration [54, 37]. Stochastic games, also known as Markov games, model the interactions of these multi-agent systems as a Markov chain over a set of states, where the transitions are triggered by joint actions and are potentially stochastic.
@@
The formalization of stochastic games was first proposed in Shapley’s seminal work [41]. A perfectly rational agent in a stochastic game is supposed to make use of all past histories to determine the next action, and therefore, the notion of strategies, defined as mappings from all possible histories to actions, is inherently complex. The fact that there are infinitely many strategies prohibits the direct application of Nash’s theorem for establishing any existence result of equilibria. However, the stationary transitions of stochastic games inevitably draw attention to a highly special subclass of time-independent and memoryless strategies that only consider the current states while discarding all past histories, termed stationary strategies. Indeed, the existence of equilibria formed by stationary strategies in $n$-player general-sum stochastic games was later proven by Fink [18] and Takahashi [48], under mild assumptions. Despite being highly restricted in terms of expressiveness, the notion of stationary strategies has enabled the community to practically investigate some complex real-world applications, particularly by resorting to multi-agent reinforcement learning (MARL) techniques, as advocated by Littman [29] and implemented in a line of subsequent work [30, 19, 39, 51].
@@
Notably, one would naturally expect strategies in other less restricted forms that can encode a broader class of behavioral patterns, hoping to achieve better payoff outcomes. For example, in the Iterated Prisoner’s Dilemma (IPD), if only stationary strategies are considered, there is a unique Nash equilibrium where both players choose Defect all the time, resulting in the lowest overall payoff. However, even with the ability to remember only one past action played by the opponent, the well-known Tit-For-Tat (TFT) strategy (having Cooperate played in the first step) can be devised. One can easily see that if both players adopt the TFT strategy, they will follow a trajectory of both playing Cooperate throughout the game, resulting in a Nash equilibrium with the highest possible payoff. Apart from other forms of representation, such as strategies represented as finite automata [40, 10, 56] and even Turing machines [33, 27, 35, 13], we focus our main effort on investigating the notion of constant-memory strategies, i.e., mappings from history segments of bounded lengths to actions, mainly because it directly relates to the concept of bounded rationality [45] in general, and is highly implementable using function approximators like Recurrent Neural Networks [24, 15] and Transformers [49] in practice.
@@
In this paper, we comprehensively study the theoretical properties associated with constant-memory strategies. We begin by presenting the following two results:
  1. 1.
@@
As a side benefit of using memories of constant length, any strategy that uses a shorter memory can always be implemented by one that uses a longer memory. Therefore, the above two results directly imply that any NE formed by shorter-length-memory strategies can be transformed into an NE formed by longer-length-memory strategies, suggesting that the longer the memory used by the strategies, the richer the equilibria one can potentially expect.
Additionally, we provide further results about best responses against mixed strategies, mathematically defined as those sampled from a set of support strategies with certain probabilities. This is associated with broad applications in the domain of opponent modeling [12, 3, 5, 55], particularly for type-based methods [3, 4, 2, 55]. However, we demonstrate that:
  1. 1.
    An Negative Result on Strategy Equivalence: An opponent with a mixed constant-memory strategy may not correspond to an equivalent opponent with a single (behavioral) constant-memory strategy in terms of achieving the same payoff.
  2. 2.
    A Negative Result on Best Responses: The best response against a mixed constant-memory strategy is not necessarily constant-memory, and computing such best responses is computationally hard.
Finally, we translate the aforementioned theoretical insights into a simple generative framework to instantiate single-agent learning environments that can be arbitrarily difficult to learn a good policy. This, in turn, reveals the potential sources of hardness underlying some single-agent decision-making problems.
2 Stochastic Games
The whole system where the agents interact is modelled as a stochastic game (SG, also known as Markov games), which can be seen as an extension of both normal-form games (to dynamic situations with stochastic transitions) and Markov decision processes (to strategic situations with multiple agents). A stochastic game is a 5-tuple $\langle\mathcal{N},\mathcal{S},\mathcal{A},T,R\rangle$ given as follows,
  1. 1.
@@ 2 Stochastic Games
Assuming complete observability and perfect recall, a perfectly rational agent should utilize the entire history, while in memory-restricted cases, an agent can only devise strategies based on past memories of finite lengths. We denote the space of all possible histories of length $K\in\mathbb{N}$ as $\mathcal{H}^{K}=(\mathcal{S}\times\mathcal{A})^{K}$. In particular, when $K=0$, we have $\mathcal{H}^{0}=\emptyset$ meaning that no history can be utilized. Then, given any non-negative integer $K$, a $K$-memory strategy for agent $i$ is a mapping from all possible histories with lengths less then or equal to $K$ and the current states to (possibly randomized) actions, mathematically denoted as $\pi_{i}:\mathcal{H}^{\leq K}\times\mathcal{S}\mapsto\Delta(\mathcal{A}_{i})$ where $\mathcal{H}^{\leq K}\triangleq\cup_{k=0}^{K}\mathcal{H}^{k}$. Let $\Pi_{i}^{K}$ denote the set of all such $K$-memory strategies for agent $i$. For convenience, we let $\mathcal{H}^{\infty}=(\mathcal{S}\times\mathcal{A})^{*}$ denote the set of complete histories that an agent with perfect recall can possibly memorize, and therefore, $\Pi_{i}^{\infty}$ is the set of all possible infinite-memory strategies for agent $i$ of the form $\pi_{i}:(\mathcal{S}\times\mathcal{A})^{*}\times\mathcal{S}\mapsto\Delta(\mathcal{A}_{i})$. A direct consequence is that $\Pi_{i}^{K}\subseteq\Pi_{i}^{K^{\prime}}\subseteq\Pi_{i}^{\infty}$ for any non-negative $K\leq K^{\prime}$. Among them, one of the most popular class of strategies is $\Pi_{i}^{0}$, termed stationary strategies. Note that an agent capable of performing infinite-memory strategies can deliberately adopt a constant-memory strategy.
@@
Note that, to compare the values of two strategy profiles, one must ensure that the limit of the right-hand side (RHS) in Equation (1) exists in the first place. Also note that, some pairs of $\pi_{i}$ and $\pi_{i}^{\prime}$ may not be comparable in the above sense, as as this comparison requires value dominance across all possible states.
3 K-Memory Best Responses and Equilibria
One should be aware of the following fact for single-agent Markov Decision Processes (MDPs) [38] in the first place, which will be considered as a lemma for the remainder of this note.
@@ 3 K-Memory Best Responses and Equilibria
  1. 1.
    Search a policy $\pi:S\mapsto\Delta(A)$ that maximizes $\mathbb{E}_{\pi}\sum_{t=0}^{\infty}[\gamma^{t}R_{t}]$, for any initial $s\in S$.
  2. 2.
    Solve the Bellman optimality equation, and then extract the policy out of the value function
    $\forall s\in S,v_{*}(s)=\max_{a\in A}\Big{[}R(s,a)+\gamma\sum_{s^{\prime}\in S}T(s^{\prime}|s,a)v_{*}(s^{\prime})\Big{]}$
@@
Given $\pi_{-i}\in\Pi_{-i}^{0}$, i.e. all other agents are adopting stationary strategies, it is sufficient for agent $i$ to best respond with a stationary strategy as well.

### v2

@@
Abstract.
Stochastic games have become a prevalent framework for studying long-term multi-agent interactions, especially in the context of multi-agent reinforcement learning. In this work, we comprehensively investigate the concept of constant-memory strategies in stochastic games. We first establish some results on best responses and Nash equilibria for behavioral constant-memory strategies, followed by a discussion on the computational hardness of best responding to mixed constant-memory strategies. Those theoretic insights are later verified on several sequential decision-making testbeds, including the Iterated Prisoner’s Dilemma, the Iterated Traveler’s Dilemma, and the Pursuit domain. This work aims to enhance the understanding of theoretical issues in single-agent planning under multi-agent systems, and uncover the connection between decision models in single-agent and multi-agent contexts. The code is available at https://github.com/Fernadoo/Const-Mem.
Key words and phrases:
@@ Abstract.
Stochastic games; Bounded rationality; Best response; Restricted memory; Reinforcement learning
1. Introduction
@@
Various real-world situations that involve long-term interactions among a group of participants can be modeled as stochastic games, such as negotiation between multiple stakeholders Baarslag et al. (2016); De Weerd et al. (2017); Baarslag (2024), bidding and mechanism design in repeated auctions Iyer et al. (2014); Balseiro et al. (2015); Guo et al. (2019); Guo et al. (2024); Su et al. (2024); Shen et al. (2020), multi-agent teamwork Sharon et al. (2015); Mirsky et al. (2022); Stern (2019), and even human-robot collaboration Zhang et al. (2023); Puig et al. (2023). Stochastic games, also known as Markov games, model the interactions of these multi-agent systems as a Markov chain over a set of states, where the transitions are triggered by joint actions and are potentially stochastic.
@@
The formalization of stochastic games was first proposed in Shapley’s seminal work Shapley (1953). A perfectly rational agent in a stochastic game is supposed to make use of all past histories to determine the next action, and therefore, the notion of strategies, defined as mappings from all possible histories to actions, is inherently complex. The fact that there are infinitely many strategies prohibits the direct application of Nash’s theorem for establishing any existence result of equilibria. However, the stationary transitions of stochastic games inevitably draw attention to a highly special subclass of time-independent and memoryless strategies that only consider the current states while discarding all past histories, termed stationary strategies. Indeed, the existence of equilibria formed by stationary strategies in $n$-player general-sum stochastic games was later proven by Fink Fink (1964) and Takahashi Takahashi (1964), under mild assumptions. Despite being highly restricted in terms of expressiveness, the notion of stationary strategies has enabled the community to practically investigate some complex real-world applications, particularly by resorting to multi-agent reinforcement learning (MARL) techniques, as advocated by Littman Littman (1994) and implemented in a line of subsequent work Lowe et al. (2017); Foerster et al. (2018); Rashid et al. (2020); Yu et al. (2022).
@@
Notably, one would naturally expect strategies in other less restricted forms that can encode a broader class of behavioral patterns, hoping to achieve better payoff outcomes. For example, in the Iterated Prisoner’s Dilemma (IPD), if only stationary strategies are considered, there is a unique Nash equilibrium where both players choose to defect all the time, resulting in the lowest overall payoff. However, even with the ability to remember only one past action played by the opponent, the well-known Tit-For-Tat (TFT) strategy (start with cooperation) can be devised. One can easily see that if both players adopt the TFT strategy, they will follow a trajectory of both cooperating throughout the game, resulting in a Nash equilibrium with the highest possible payoff. Apart from other forms of representation, such as strategies represented as finite automata Rubinstein (1986); Ben-Porath (1990); Zuo and Tang (2015) and even Turing machines Megiddo and Wigderson (1986); Knoblauch (1994); Nachbar and Zame (1996); Chen et al. (2017b), we focus our main effort on investigating the notion of constant-memory strategies, i.e., mappings from history segments of bounded lengths to actions, mainly because it directly relates to the concept of bounded rationality Simon (1990) in general, and is highly implementable using function approximators like Recurrent Neural Networks Hochreiter and Schmidhuber (1997); Cho et al. (2014) and Transformers Vaswani et al. (2017) in practice. Note that this notion has been preliminarily investigated by Chen et al. Chen et al. (2017a) and Wang and Lin Wang and Lin (2019). However, they only focus on behavioral strategy best responses for repeated games, without further discussion on either Nash equilibria or mixed strategies.
@@
In this paper, we comprehensively study the theoretical properties associated with constant-memory strategies in stochastic games. We begin by presenting the following two results:
  1. (1)
@@
As a side benefit of using memories of constant lengths, any strategy that uses a shorter memory can always be implemented by one that uses a longer memory. Therefore, the above two results directly imply that any NE formed by shorter-length-memory strategies can be transformed into an NE formed by longer-length-memory strategies, suggesting that the longer the memory used by the strategies, the richer the equilibria one can potentially expect.
Additionally, we provide further results about best responses against mixed constant-memory strategies, mathematically defined as those sampled from a set of support strategies with certain probabilities. This is associated with broad applications in the domain of opponent modeling Carmel and Markovitch (1998); Albrecht and Ramamoorthy (2015); Albrecht and Stone (2018); Zhu and Lin (2025), particularly for type-based methods Albrecht and Ramamoorthy (2015, 2019); Albrecht et al. (2016); Zhu and Lin (2025). However, we demonstrate that:
  1. (1)
    An Negative Result on Strategy Equivalence: An opponent with a mixed constant-memory strategy may not correspond to an equivalent opponent with a single (behavioral) constant-memory strategy in terms of resulting in the same payoff.
  2. (2)
    A Negative Result on Best Responses: The best response against a mixed constant-memory strategy is not necessarily constant-memory, and computing such best responses is computationally hard, possibly even not computable.
In spite of these negative results, we do provide a computational model for solving the best response against a mixed strategy, which 1) also serves as the evidence that those computational models proposed by Zhu and Lin Zhu and Lin (2025) do not over-complicate the problem; and 2) can be carried over to the methods in Zhu and Lin (2025) which only assume stationary strategies.
2. Preliminaries
The whole system where the agents interact is modelled as a stochastic game (SG, also known as Markov games) (Shapley, 1953; Solan and Vieille, 2015), which can be seen as an extension of both normal-form games (to dynamic situations with stochastic transitions) and Markov decision processes (to strategic situations with multiple agents). A stochastic game is a 5-tuple $\langle\mathcal{N},\mathcal{S},\mathcal{A},T,R\rangle$ given as follows,
  1. (1)
@@ 2. Preliminaries
Assuming complete observability and perfect recall, a perfectly rational agent should utilize the entire history, while in memory-restricted cases, an agent can only devise strategies based on past memories of finite lengths. We denote the space of all possible histories of length $K\in\mathbb{N}$ as $\mathcal{H}^{K}\triangleq(\mathcal{S}\times\mathcal{A})^{K}$. In particular, when $K=0$, we have $\mathcal{H}^{0}=\emptyset$ meaning that no history can be utilized. Then, given any non-negative integer $K$, a $K$-memory strategy for agent $i$ is a mapping from all possible histories with lengths less then or equal to $K$ and the current states to (possibly randomized) actions, mathematically denoted as $\pi_{i}:\mathcal{H}^{\leq K}\times\mathcal{S}\mapsto\Delta(\mathcal{A}_{i})$ where $\mathcal{H}^{\leq K}\triangleq\cup_{k=0}^{K}\mathcal{H}^{k}$. Let $\Pi_{i}^{K}$ denote the set of all such $K$-memory strategies for agent $i$. For convenience, we let $\mathcal{H}^{\infty}=(\mathcal{S}\times\mathcal{A})^{*}$ denote the set of complete histories that an agent with perfect recall can possibly memorize, and therefore, $\Pi_{i}^{\infty}$ is the set of all possible infinite-memory strategies for agent $i$ of the form $\pi_{i}:(\mathcal{S}\times\mathcal{A})^{*}\times\mathcal{S}\mapsto\Delta(\mathcal{A}_{i})$. A direct consequence is that $\Pi_{i}^{K}\subseteq\Pi_{i}^{K^{\prime}}\subseteq\Pi_{i}^{\infty}$ for any non-negative $K\leq K^{\prime}$. Among them, one of the most popular class of strategies is $\Pi_{i}^{0}$, termed stationary strategies. Note that an agent capable of performing infinite-memory strategies can deliberately adopt a constant-memory strategy. To be clear, we use the term constant-memory to distinguish from those infinite-memory strategies, and use the term K-memory when this specific $K$ needs to be emphasized.
@@
requiring that a $\pi_{i}$ must outperform any other in $\Pi_{i}^{\infty}$ to serve as the best response. Note that, to compare the values of two strategy profiles, one must ensure that the limit of the right-hand side (RHS) in Equation (1) exists in the first place. Also note that, some pairs of $\pi_{i}$ and $\pi_{i}^{\prime}$ may not be comparable in the above sense, as as this comparison requires value dominance across all possible states.
3. Best Responses and Nash Equilibria
One should be aware of the following fact for single-agent Markov Decision Processes (MDPs) Puterman (2014) in the first place, which will be considered as a lemma for the remainder of this paper.
@@ 3. Best Responses and Nash Equilibria
  1. (1)
    Searching for a policy $\pi_{*}:S\mapsto\Delta(A)$ that maximizes the accumulated rewards $\mathbb{E}_{\pi_{*}}\sum_{t=0}^{\infty}[\gamma^{t}R_{t}]$, for any initial $s\in S$.
  2. (2)
    Solving the Bellman optimality equation below
    $\forall s\in S,v_{*}(s)=\max_{a\in A}\Big[R(s,a)+\gamma\sum_{s^{\prime}\in S}T(s^{\prime}|s,a)v_{*}(s^{\prime})\Big],$
    and then extracting the policy from the optimal value function
    $\pi_{*}(s)\in\arg\max_{a\in A}\Big[R(s,a)+\gamma\sum_{s^{\prime}\in S}T(s^{\prime}|s,a)v_{*}(s^{\prime})\Big].$
@@
We first characterize the best response of an agent when all the other opponents are equipped with constant-memory strategies with the same non-negative (and finite) memory length.
@@
Given $\pi_{j}\in\Pi_{j}^{K}$ with $K\in\mathbb{Z}$ for all $j\neq i$ , i.e., all the other agents are adopting constant-memory strategies with the same finite memory length $K$, it is sufficient for agent $i$ to best respond with a $K$-memory strategy as well.
Proof sketch.
As the full proof involves some fundamental (and probably tedious) derivations, we defer it to Appendix A.1. The main issue is that, in general the decision process from the perspective of agent $i$ is an infinite MDP where states encompass infinitely many histories (of all possible lengths), and transitions/rewards are jointly controlled by the stochastic game itself as well as the other opponents. Thus, the goal of this proof is to show that there indeed exists a finite MDP with the same effect. Here, we present a proof sketch by construction, which is, in fact, a consequence of the full proof, and can be approached from a more direct perspective.
Given $\pi_{j}\in\Pi_{j}^{K}$ for all $j\neq i$, agent $i$ is then faced with an MDP with environmental states augmented by finite-length histories, denoted as $\mathcal{M}^{K}(\pi_{-i})=\langle\mathcal{H}^{\leq K}\times\mathcal{S},\mathcal{A}_{i},T^{K}_{\pi_{-i}},R^{K}_{\pi_{-i}},\gamma\rangle$,
  • $\mathcal{A}_{i}$ and $\gamma$ are inherited from the previous setup,
  • A state is now consisting of the past $K$ steps plus the current environmental state, resulting in a space of $\mathcal{H}^{\leq K}\times\mathcal{S}$,
  • Transitions are now made among the augmented states, i.e., for every pair $(H^{\prime},S^{\prime})$, $(H,S)\in\mathcal{H}^{\leq K}\times\mathcal{S}$, and $a_{i}\in\mathcal{A}_{i}$,
[…]
  • The reward for each $(H,S)\in\mathcal{H}^{\leq K}\times\mathcal{S}$ and $a_{i}\in\mathcal{A}_{i}$ is
    $R^{K}_{\pi_{-i}}(H,S,a_{i})\triangleq\sum_{a_{-i}\in\mathcal{A}_{-i}}R_{i}(S,a)\pi_{-i}(a_{-i}|H,S)$
Among all the optimal solutions of $\mathcal{M}^{K}(\pi_{-i})$, there must exist a stationary (and deterministic) policy, i.e. $\pi_{*}:\mathcal{H}^{\leq K}\times\mathcal{S}\mapsto\mathcal{A}_{i}$, which corresponds to a $K$-memory (and pure) strategy best response of agent $i$ against $\pi_{-i}\in\Pi_{-i}^{K}$.
This proof is summarized in our code implementation. [^1] Please refer to the notebook code/kMemBR.ipynb in the codebase.
One can immediately see the following corollary where the opponents may use constant-memory strategies but with potentially different memory lengths. The justification is straightforward: all the opponents can be jointly viewed as a “super-agent”, and consequently this “super-agent” is adopting a $(\max\{K_{j}\}_{j\neq i})$-memory strategy. Therefore, the best response for the pivotal agent shall also be of $(\max\{K_{j}\}_{j\neq i})$-memory.
Corollary 0.
Given $\pi_{j}\in\Pi_{j}^{K_{j}}$ with each $K_{j}\in\mathbb{Z}$ for all $j\neq i$, i.e., all the other agents are adopting constant-memory strategies but with varying memory lengths, it is sufficient for agent $i$ to best respond with a $(\max\{K_{j}\}_{j\neq i})$-memory strategy.
As best responses are well established, we will examine whether an equilibrium exists when everyone best responds to one another.
Definition 0 (Nash Equilibrium).
$\forall i\in\mathcal{N},\pi_{i}^{*}\in BR(\pi_{-i}^{*})$
Lemma 0 (Brouwer’s Fixed-Point Theorem Brouwer (1911)).
Let $\Delta=\prod_{l=1}^{L}\Delta_{m_{l}}$, where each $\Delta_{m_{l}}$ is a simplex in $\mathbb{R}^{m_{l}+1}$. If $f:\Delta\mapsto\Delta$ is a continuous mapping, then $f$ has a fixed point.
Theorem 6.
There exists an NE when the agents are all adopting $K$-memory strategies, for any arbitrary non-negative finite $K$.